# Hybrid quantum cycle generative adversarial network for small molecule generation

### v1

@@ Abstract
The contemporary drug design process demands considerable time and resources to develop each new compound entering the market. Generating small molecules is a pivotal aspect of drug discovery, essential for developing innovative pharmaceuticals. Uniqueness, validity, diversity, druglikeliness, synthesizability, and solubility molecular pharmacokinetic properties, however, are yet to be maximized. This work introduces several new generative adversarial network models based on engineering integration of parametrized quantum circuits into known molecular generative adversarial networks. The introduced machine learning models incorporate a new multi-parameter reward function grounded in reinforcement learning principles. Through extensive experimentation on benchmark drug design datasets, QM9 and PC9, the introduced models are shown to outperform scores achieved previously. Most prominently, the new scores indicate an increase of up to $30\%$ in the druglikeness quantitative estimation. The new hybrid quantum machine learning algorithms, as well as the achieved scores of pharmacokinetic properties, contribute to the development of fast and accurate drug discovery processes.
@@ I Introduction
In the current pharmaceutical landscape, the drug design process is a prolonged and costly endeavor. It typically spans up to $15$ years Hughes_Rees_Kalindjian_Philpott_2011 from target identification to clinical application, incurring expenses of approximately $1$ billion for each new drug. Machine Learning has shown successful uses through the different stages of drug development, from the search for specific protein inhibitors Wang_Chen_Yang_Akutsu_2022 to the evaluation of pharmacokinetic properties and adverse effects.
Generative Adversarial Networks (GANs) goodfellow2014generative have gained prominence in molecular design. Their architecture is adept at generating a vast array of potential drug candidates from extensive molecular spaces, thereby facilitating more efficient preliminary screenings. GAN models, especially when compared to recurrent neural networks schmidt2019recurrent and variational autoencoders Kingma_2019 , have demonstrated superiority in generating SMILES Weininger_1988 representations of compounds. The use of graph representations of compounds, which are invariant to the permutations of atom orders, has allowed GANs, particularly the MolGAN decao2022molgan , to become state-of-the-art approach in generative chemistry.
@@
Utilization of hybrid quantum neural networks (HQNNs) represents a convergence of classical deep learning architectures with Quantum Machine Learning (QML) algorithms jerbi2023quantum ; perez2022reduce ; marshall2022high ; kordzanganeh2023parallel , specifically through Parameterized Quantum Circuits (PQCs) senokosov2023quantum . This hybrid approach harnesses the strengths of classical and quantum computing, introducing a system capable of efficiently processing large datasets compared to classical deep learning architectures alone li2020quantum ; Mitarai2020 . HQNNs have exhibited promising applications across various industrial domains, including healthcare houssein2022hybrid ; lusnig2023hybrid ; jain2022hybrid , chemistry sedykh2023quantum ; kurkin2023forecasting , routing haboury2023supervised and aerospace rainjonneau2023quantum , also in image classification sagingalieva2023hyperparameter ; landman2022quantum . While HQNNs have demonstrated efficacy in these fields, further research is essential to explore their potential in drug design fully and to refine algorithms for enhanced robustness and scalability.
This article introduces the Hybrid Quantum Cycle MolGAN for generating graph representations of small molecules. By incorporating a Cycle Component into the Hybrid Quantum MolGAN (HQ-MolGAN), where both the Generator and the Cycle Component are represented using an HQNN hybridTQ2022 ; qml_review_2023 , we have been able to stabilize the training process for molecular samples and improve key metrics. This includes increases of up to $30\%$ in the Quantitative Estimate of Druglikeness (QED score) Bickerton_Paolini_Besnard_Muresan_Hopkins_2012 . Such models, have shown quantum superiority over classical MolGAN in all three key metrics QED, Synthesis Accessibility score (SA) Ertl_Schuffenhauer_2009 and logP (characteristic of the hydrophilicity and lipophilicity of a drug) Leo_Hansch_Elkins_1971 ), as well as an advantage over nearest competitors among quantum models.
@@ II Results
In this study, we introduce several models for small molecule generation. Firstly, we present a refined classical MolGAN with a halved parameter count, drawing inspiration from the state-of-the-art classical MolGAN architecture decao2022molgan . Secondly, in Sec. II we present an HQ-MolGAN, fusing classical and quantum computing approaches for enhanced capabilities. Notably, we propose the novel classical Cycle MolGAN, incorporating a multi-parameter reward function based on reinforcement learning principles, inspired by the state-of-the-art Cycle MolGAN Maziarka_Pocha_Kaczmarczyk_Rataj_Danel_Warchoł_2020a . Additionally, in Sec. II, we introduce the innovative Hybrid Quantum Cycle MolGAN (HQ-Cycle-MolGAN). Through rigorous experimentation on the QM9 and PC9 datasets, described in Sec. II, our results demonstrate that models trained on PC9 exhibit higher LogP scores than their QM9 counterparts. Furthermore, the hybrid quantum models showcase superior performance, achieving the highest QED, SA, and LogP metrics scores. Remarkably, our hybrid quantum model outperforms a similar QGAN-HG MR hybrid model from Ref. Li_Topaloglu_Ghosh_2021 and QuMolGAN from Ref. Kao_2023 , emphasizing the efficacy of our proposed approaches in small molecule generation. We summarize our conclusions and outline future research directions in Sec. III.
@@
This study employed two datasets for model training: QM9 and PC9. In this paper, two datasets were used for the training procedure. The QM9 dataset ramakrishnan2014quantum ; Ruddigkeit_vanDeursen_Blum_Reymond_2012 , a well-established benchmark in small molecule drug design since 2012, comprises approximately $134,000$ neutral molecules, each with no more than nine atoms $(C,O,N,F)$ apart from hydrogen. Its comprehensive and diverse chemical space makes it particularly relevant for this field.
@@
Fig. 1 shows that the mean QED and LogP scores for molecules in the PC9 dataset are higher than those in QM9, while the mean score for Synthesis Accessibility (SA) is lower. Theoretically, this suggests that models trained on the PC9 dataset might be inclined to generate samples with higher values in these two key metrics than similar models trained on QM9. However, as detailed in Sec. II, this is not always the case.
@@
Figure 2: (a) Structure of HQ-Cycle-MolGAN: Generator (G), Discriminator (D), Cycle Component (C). Highlighted with green part of HQ-Cycle-MolGAN is same for HQ-MolGAN. (b) Illustration of work of Cycle Components Let’s suppose Z is a space of normally distributed noise vectors, $Y$-chemical space of datasets. Then Generator maps $X$ to some chemical space $Y^{\prime}$, and after Cycle Component restores vector G(Z) back to noise. Then accuracy of these restorations is optimized. (c) Quantum Depth-Infused Neural Network Layer that was used as HQ-Cycle component.
Figure 3: Chart of (a) QED, (b) LogP, and (c) SA scores during training of classical MolGAN and Hybrid MolGAN. It can be seen that while MolGAN limits its scores to a narrow beam of values even after $50,000$ iterations, when Hybrid MolGAN presents a wider range of compounds, covering greater scores of key metrics.
In this section we introduce the HQ-MolGAN, which is based on classical MolGAN architecture decao2022molgan . As depicted in Fig.2 (a), green rectangular, the architecture of HQ-MolGAN is comprised of three primary components: the Generator (G), the Discriminator (D), and the Reward component (R). This model operates on the principles of the Wasserstein Generative Adversarial Network arjovsky2017wasserstein , wherein the Generator endeavors to synthesize molecular graph representations indistinguishable from authentic ones, thereby “deceiving” the Discriminator. The training regimen of HQ-MolGAN encapsulates a min-max optimization game, wherein the Generator and Discriminator engage in a continuous adaptive process to refine the generative quality of molecular representations:
@@
The Reward Component in our HQ-MolGAN architecture functions as a sophisticated Reinforcement Learning objective, tasked with evaluating the Generator’s output based on several chemical property metrics. This evaluation extends beyond the conventional metrics of Quantum Estimation of QED, LogP, and SA scores. It incorporates a comprehensive assessment of “validity”, quantified as the ratio of valid molecular samples to the total number of generated samples. Furthermore, it considers “novelty”, defined by the proportion of generated valid samples not present in the training dataset. Additionally, the Reward Component assesses “diversity”, a measure of the variance in the chemical structures of the generated molecules, and the “Natural Product likeness” (NP) score, as delineated by Ertl_Roggo_Schuffenhauer_2007 . These multifaceted evaluation criteria enable a more nuanced and thorough assessment of the Generator’s performance, aligning the generated molecules more closely with desired chemical characteristics.
In the architecture of HQ-MolGAN, a pivotal role is played by the Variational Quantum Circuit (VQC), which is integrated as the initial layer in MolGAN’s generator. The VQC operates by encoding a noise vector into $N$ qubits. Subsequent to the application of rotation and entanglement gates, the VQC outputs a probability distribution vector, denoted as $[p(0),...,p(2^{N}-1)]$, where each element represents the probability of a corresponding quantum state. This vector, with a dimensionality of $2^{N}$, undergoes a truncation process where only the first $2^{N-N_{\text{ancilla}}}$ elements are retained. The truncated vector is then fed into a classical fully-connected layers, which constitutes the remaining component of HQ-MolGAN’s generator.
@@
The operational mechanism of the VVRQ layer within our HQ-MolGAN framework (Fig. 2 (a)) is as follows: encoding the noise vector directly onto the qubits in a single step using angle embedding PhysRev.70.460 (green rectangles in Fig. 2 (a)). Following this initialization, the VVRQ layer implements several variational layers which consist of a sequence of Rot (rotation) gates (blue rectangles in Fig. 2 (a))
@@
$Ry(\theta)=\begin{pmatrix}\text{cos}(\frac{\theta}{2})&-\text{sin}(\frac{\theta}{2})\\ \text{sin}(\frac{\theta}{2})&\ \text{cos}(\frac{\theta}{2})\end{pmatrix}\ \ Rz(\theta)=\begin{pmatrix}\text{exp}(-\frac{\theta}{2})&0\\ 0&\text{exp}(\frac{\theta}{2})\end{pmatrix}$
@@
$\text{CNOT}=\begin{pmatrix}1&0&0&0\\ 0&0&0&1\\ 0&0&1&0\\ 0&1&0&0\\ \end{pmatrix}$
Originally this approach was proposed in Tsang_West_Erfani_Usman_2023 for image generation, but it can be applied to any generative task. The CNOT gates are applied between sequentially adjacent qubits, i.e., between qubit $i$ and qubit $i+1$, thereby facilitating quantum entanglement and information propagation across the qubit array.
@@
In the EFQ layer the data encoding process is distinctly characterized by a dual-phase approach. Initially, the input data is encoded onto the qubits using angle embedding. This is followed by several variational layers, and after this, a second encoding phase, wherein the amplitude of the rotational gates is systematically increased to double its initial value.
This dual encoding scheme, particularly with the amplified rotational amplitude in the second phase, is designed to enhance the expressive power of the quantum circuit schuld_fourier . By manipulating the amplitude of rotations in this manner, the EFQ layer is potentially able to induce a more diverse and complex quantum state space.
@@
The Cycle-MolGAN model introduces an innovative “Cycle component” (C) (Fig. 2 (b)) to the established MolGAN architecture. This component is ingeniously designed to reverse the molecule generation process. Specifically, it converts the generated molecular samples back into their originating noise vectors and assesses the accuracy of this reverse conversion. This approach, as proposed by Maziarka_Pocha_Kaczmarczyk_Rataj_Danel_Warchoł_2020a , has been identified as particularly advantageous in the realm of molecular optimization tasks. It contributes significantly to the stability of the training performance and is instrumental in suppressing the training of the non-isomorphic generator compounds within the Hybrid-MolGAN framework, especially pertinent in the generation of small molecules.
@@
$L(\theta)=\lambda\cdot L(\theta)_{\text{WGAN}}+(1-\lambda)\cdot L(\theta)_{\text{Cycle}}+\gamma L(\theta)_{\text{Reward}}\\ \gamma,\ \lambda\in[0,1]\end{aligned}.$
@@
Figure 4: (a) Samples generated by HQ-MolGAN-VVRQ trained on QM9. (b) “Highly entropy state”: HQ-MolGAN-VVRQ generated inappropriate samples and RDKit rewarded them with average metric of LogP $\propto$ 0.9. (c) Samples generated by HQ-Cycle-MolGAN-VVRQ trained on Both datasets. (d) Samples generated by MolGAN with HQ-Cycle trained on Both datasets.
The models in this study were implemented using Python3, leveraging the PyTorch framework PyTorch and PennyLane Pennylane for quantum computations. For the evaluation of the chemical properties of the synthesized compounds, the RDKit library was employed. The computational experiments were conducted on high-performance GPU hardware, specifically the Tesla V100 and the RTX 3060.
For the classical MolGAN models, the generator’s architecture was scaled down by reducing the number of parameters in each layer by half, resulting in a total of $157,570$ parameters in the Generator. The classical models, including both the standard MolGAN and Cycle-MolGAN, underwent a training regime of $200,000$ iterations with a batch size of $10$ samples. In contrast, the hybrid quantum models were subjected to a shorter training duration of $50,000$ iterations. The validation set size was limited, containing either $100$ samples when training on a single dataset or $250$ samples in cases where multiple datasets were employed. In both EFQ and VVRQ models number of ancilia qubits is equal to $2$.
The experimental investigation was conducted in three distinct stages:
@@
These three stages are described in detail in the next three subsections.

### v2

@@ Abstract
The drug design process currently requires considerable time and resources to develop each new compound that enters the market. This work develops an application of hybrid quantum generative models based on the integration of parametrized quantum circuits into known molecular generative adversarial networks, and proposes quantum cycle architectures that improve model performance and stability during training. Through extensive experimentation on benchmark drug design datasets, QM9 and PC9, the introduced models are shown to outperform the previously achieved scores. Most prominently, the new scores indicate an increase of up to $30\%$ in the quantitative estimation of druglikeness. The new hybrid quantum machine learning algorithms, as well as the achieved scores of pharmacokinetic properties, contribute to the development of fast and accurate drug discovery processes.
@@ I Introduction
In the current pharmaceutical landscape, the drug design process is a prolonged and costly endeavor. It typically spans up to $15$ years Hughes_Rees_Kalindjian_Philpott_2011 from target identification to clinical application, incurring expenses of approximately \$$\$1$ billion for each new drug. Machine learning has shown successful uses through the different stages of drug development, from the search for specific protein inhibitors Wang_Chen_Yang_Akutsu_2022 to the evaluation of pharmacokinetic properties and adverse effects.
Generative Adversarial Networks (GANs) goodfellow2014generative have gained prominence in molecular design. Their architecture is adept at generating a vast array of potential drug candidates from extensive molecular spaces, thereby facilitating more efficient preliminary screenings. GAN models, especially when compared to recurrent neural networks schmidt2019recurrent and variational autoencoders Kingma_2019 , have demonstrated superiority in generating SMILES Weininger_1988 representations of compounds. A novel quantum approach introduced in Ref. Gircha_Boev_Avchaciov_Fedichev_Fedorov_2023 used the Deep Variational Autoencoder model trained to construct molecules as SMILES strings. The advancements in molecule representations led to the use of graph representations of compounds. The use of graphs instead of SMILES, which are invariant to the permutation of atom orders, has allowed GANs, particularly MolGAN decao2022molgan , to become the state-of-the-art approach in generative chemistry.
@@
The study of hybrid quantum neural networks (HQNNs) represents a convergence of classical deep learning architectures with quantum machine learning (QML) algorithms qml_review_2023 ; jerbi2023quantum ; perez2022shallow ; marshall2023highdimensional ; kordzanganeh2023parallel , specifically through parameterized quantum circuits senokosov2024quantum . This hybrid approach harnesses the strengths of classical and quantum computing, introducing a system capable of efficiently processing large datasets compared to classical deep learning architectures alone li2020quantum ; Mitarai2020 . HQNNs have exhibited promising applications across various industrial domains, including healthcare houssein2022hybrid ; lusnig2024hybrid ; jain2022hybrid , chemistry sedykh2024hybrid ; kurkin2023forecasting , the energy industry sagingalieva2023photovoltaic , routing haboury2023supervised and aerospace rainjonneau2023quantum , as well as in image classification sagingalieva2023hyperparameter ; landman2022quantum . While HQNNs have demonstrated efficacy in these fields, further research is essential to explore their potential in drug design.
This article introduces the Hybrid Quantum Cycle MolGAN for generating graph representations of small molecules. By incorporating a Cycle Component into the Hybrid Quantum MolGAN (HQ-MolGAN), where both the Generator and the Cycle Component are represented using an HQNN hybridTQ2022 ; qml_review_2023 , we have been able to stabilize the training process for molecular samples and improve key metrics. This includes increases of up to $30\%$ in the Quantitative Estimate of Druglikeness (QED score) Bickerton_Paolini_Besnard_Muresan_Hopkins_2012 , a composite metric that evaluates a molecule’s overall drug-likeness based on its chemical structure. The QED score is instrumental in assessing the potential of a compound to qualify as an effective drug, providing a quantitative measure that can guide early drug discovery efforts. Additionally, we have observed improvements in the Synthesis Accessibility score (SA) Ertl_Schuffenhauer_2009 and the logP score Leo_Hansch_Elkins_1971 . The SA score quantifies the complexity of synthesizing a given molecular compound, offering insights into the practicality of its production at scale. A lower SA score indicates easier synthesis, which is favorable for drug development. The LogP score measures a compound’s solubility and permeability, indicating its balance between hydrophilicity (water solubility) and lipophilicity (fat solubility). This balance is crucial for a drug’s absorption, distribution, metabolism, and excretion properties, impacting its effectiveness and safety. Overall, the proposed hybrid models show an advantage over their nearest competitors among quantum models in terms of pharmacokinetic properties, including QED score, SA score, and logP score.
@@ II Results
In this study, we introduce several models for small molecule generation. Firstly, we present a refined classical MolGAN with a halved parameter count, drawing inspiration from the state-of-the-art classical MolGAN architecture decao2022molgan . Secondly, in Sec. II, we present an HQ-MolGAN, fusing classical and quantum computing approaches for enhanced capabilities. Notably, we propose the novel classical Cycle MolGAN, incorporating a multi-parameter reward function based on reinforcement learning principles, inspired by the state-of-the-art Cycle MolGAN Maziarka_Pocha_Kaczmarczyk_Rataj_Danel_Warchoł_2020a . Additionally, in Sec. II, we introduce the innovative Hybrid Quantum Cycle MolGAN (HQ-Cycle-MolGAN). Through rigorous experimentation on the QM9 and PC9 datasets, described in Sec. II, our results demonstrate that models trained on PC9 exhibit higher LogP scores than their QM9 counterparts. Furthermore, the hybrid quantum models showcase better performance, achieving the highest QED, SA, and LogP metrics scores. Remarkably, our hybrid quantum model outperforms a similar QGAN-HG MR hybrid model from Ref. Li_Topaloglu_Ghosh_2021 and QuMolGAN from Ref. Kao_2023 , emphasizing the efficacy of our proposed approaches in small molecule generation. We summarize our conclusions and outline future research directions in Sec. III.
@@
This study employed two datasets for model training: QM9 and PC9. The QM9 dataset ramakrishnan2014quantum ; Ruddigkeit_vanDeursen_Blum_Reymond_2012 , a well-established benchmark in small molecule drug design since 2012, comprises approximately $134,000$ neutral molecules, each with no more than nine atoms $(C,O,N,F)$ apart from hydrogen. Its comprehensive and diverse chemical space makes it particularly relevant for this field.
@@
Fig. 1 shows that the mean QED and LogP scores for molecules in the PC9 dataset are higher than those in QM9, while the mean score for Synthesis Accessibility (SA) is lower. Intuitively, this suggests that models trained on the PC9 dataset might be inclined to generate samples with higher values in these two key metrics than similar models trained on QM9. However, as detailed in Sec. II, this is not always the case.
@@
Figure 2: (a) Structure of HQ-Cycle-MolGAN: Generator (G), Discriminator (D), Cycle Component (C). The part highlighted in green is the same as HQ-MolGAN. (b) Illustration of the work of the Cycle Components. Suppose $Z$ is a space of normally distributed noise vectors, and $Y$ is a chemical space of datasets. The Generator maps $Z$ to some chemical space $Y^{\prime}$, and after the Cycle Component restores vector $G(Z)$ back to noise. The accuracy of these restorations is then optimized. (c) Quantum Depth-Infused Neural Network Layer used as the HQ-Cycle component.
Figure 3: Chart of (a) QED, (b) LogP, and (c) SA scores during the training of classical MolGAN and Hybrid MolGAN. It can be seen that while MolGAN limits its scores to a narrow beam of values even after $50,000$ iterations, Hybrid MolGAN presents a wider range of compounds, covering greater scores of key metrics.
In this section, we introduce the HQ-MolGAN, which is based on the classical MolGAN architecture decao2022molgan . As depicted in the shaded green rectangle of Fig. 2(a), the architecture of HQ-MolGAN is comprised of three primary components: the Generator (G), the Discriminator (D), and the Reward component (R). This model operates on the principles of the Wasserstein Generative Adversarial Network arjovsky2017wasserstein , wherein the Generator endeavors to synthesize molecular graph representations indistinguishable from authentic ones, thereby “deceiving” the Discriminator. The training regimen of HQ-MolGAN encapsulates a min-max optimization game, wherein the Generator and Discriminator engage in a continuous adaptive process to refine the generative quality of molecular representations:
@@
The Reward Component in our HQ-MolGAN architecture functions as a sophisticated Reinforcement Learning objective, tasked with evaluating the Generator’s output based on several chemical property metrics. This evaluation extends beyond the conventional metrics of QED, LogP, and SA scores. It incorporates a comprehensive assessment of “validity,” quantified as the ratio of valid molecular samples to the total number of generated samples. Furthermore, it considers “novelty,” defined by the proportion of generated valid samples not present in the training dataset. Additionally, the Reward Component assesses “diversity,” a measure of the variance in the chemical structures of the generated molecules, and the “Natural Product likeness” (NP) score, as delineated by Ref. Ertl_Roggo_Schuffenhauer_2007 . These multifaceted evaluation criteria enable a more nuanced and thorough assessment of the Generator’s performance, aligning the generated molecules more closely with desired chemical characteristics.
In the architecture of HQ-MolGAN, a pivotal role is played by the Variational Quantum Circuit (VQC), which is integrated as the initial layer in MolGAN’s generator. The VQC operates by encoding a noise vector into $N$ qubits. Subsequent to the application of rotation and entanglement gates, the VQC outputs a probability distribution vector, denoted as $[p(0),...,p(2^{N}-1)]$, where each element represents the probability of a corresponding quantum state. This vector, with a dimensionality of $2^{N}$, undergoes a truncation process where only the first $2^{N-N_{\text{ancilla}}}$ elements are retained. The truncated vector is then fed into the classical fully-connected layers, which constitute the remaining component of HQ-MolGAN’s generator.
@@
The operational mechanism of the VVRQ layer within our HQ-MolGAN framework (Fig. 2(a)) involves encoding the noise vector directly onto the qubits in a single step using angle embedding PhysRev.70.460 (green rectangles in Fig. 2(a)). Following this initialization, the VVRQ layer implements several variational layers which consist of a sequence of Rot (rotation) gates (blue rectangles in Fig. 2(a))
@@
$Ry(\theta)=\begin{pmatrix}\cos(\frac{\theta}{2})&-\sin(\frac{\theta}{2})\\[4.0pt] \sin(\frac{\theta}{2})&\ \cos(\frac{\theta}{2})\end{pmatrix}$
$Rz(\theta)=\begin{pmatrix}\text{exp}(-i\frac{\theta}{2})&0\\ 0&\text{exp}(i\frac{\theta}{2})\end{pmatrix}$
@@
$\text{CNOT}=\begin{pmatrix}1&0&0&0\\ 0&1&0&0\\ 0&0&0&1\\ 0&0&1&0\\ \end{pmatrix}$
Originally this approach was proposed in Ref. Tsang_West_Erfani_Usman_2023 for image generation, but it can be applied to any generative task. The CNOT gates are applied between sequentially adjacent qubits, i.e., between qubit $i$ and qubit $i+1$, thereby facilitating quantum entanglement and information propagation across the qubit array.
@@
In the EFQ layer, the data encoding process is distinctly characterized by a dual-phase approach. Initially, the input data is encoded onto the qubits using angle embedding. This is followed by several variational layers and a second encoding phase, in which the amplitude of the rotational gates is systematically increased to double its initial value.
This dual encoding scheme, particularly with the amplified rotational amplitude in the second phase, is designed to enhance the expressive power of the quantum circuit schuld_fourier . By manipulating the amplitude of rotations in this manner, the EFQ layer could potentially induce a more diverse and complex quantum state space.
@@
The Cycle-MolGAN model introduces an innovative “Cycle component” (C) (Fig. 2(b)) to the established MolGAN architecture. This component is ingeniously designed to reverse the molecule generation process. Specifically, it converts the generated molecular samples back into their originating noise vectors and assesses the accuracy of this reverse conversion. This approach, as proposed by Ref. Maziarka_Pocha_Kaczmarczyk_Rataj_Danel_Warchoł_2020a ; 8237506 , has been identified as particularly advantageous in the realm of molecular optimization tasks. It contributes significantly to the stability of the training performance and is instrumental in suppressing the training of the non-isomorphic generator compounds within the Hybrid-MolGAN framework, especially pertinent in generating small molecules.
@@
$L(\theta)$ $=\lambda\cdot L(\theta)_{\text{WGAN}}+(1-\lambda)\cdot L(\theta)_{\text{Cycle}}+\gamma L(\theta)_{\text{Reward}}$
$\gamma,\ \lambda$ $\in[0,1].$
@@
Figure 4: (a) Samples generated by HQ-MolGAN-VVRQ trained on QM9. (b) “High entropy state”: HQ-MolGAN-VVRQ generated inappropriate samples and RDKit rewarded them with an average metric of LogP $\propto$ 0.9. (c) Samples generated by HQ-Cycle-MolGAN-VVRQ trained on both datasets. (d) Samples generated by MolGAN with HQ-Cycle trained on both datasets.
The models in this study were developed in Python3, utilizing the PyTorch framework PyTorch and PennyLane Pennylane for the quantum computation. The simulations described further were performed on classical simulators emulating quantum hardware. To evaluate the chemical properties of the synthesized compounds, we utilized the RDKit library. Our computational experiments leveraged GPU hardware, specifically the Tesla V100 and RTX 3060 GPUs, to facilitate efficient processing.
For testing the potential performance on quantum hardware, a Qiskit Qiskit implementation of the VQC was used on the simulator of the IBM Brisbane device IBMsimulator1 ; IBMsimulator2 .
For the classical MolGAN models, the generator’s architecture was scaled down by reducing the number of parameters in each layer by half, resulting in a total of $157,570$ parameters in the Generator. The classical models, including both the standard MolGAN and Cycle-MolGAN, underwent a training regime of $200,000$ iterations with a batch size of $10$ samples. In contrast, the hybrid quantum models were subjected to a shorter training duration of $50,000$ iterations. The validation set size was limited, containing either $100$ samples when training on a single dataset or $250$ samples in cases where multiple datasets were employed. In both the EFQ and VVRQ models, the number of ancilla qubits is equal to $2$.
The experimental investigation was conducted in four distinct stages:
@@
  • The Fourth Stage (Sec. II.4): This phase included a setup and analysis of HQ-MolGAN’s performance test after forward pass on the classical simulator of the IBM Brisbane quantum device.
These four stages are described in detail in the next three subsections.